OpenStudy (kicker):

You travel 400 km in 4 h. Your _____ speed is 100 km/h. constant average instantaneous initial

OpenStudy (anonymous):

constant, because you don't change

OpenStudy (auctoratrox):

the answer is average because over the entire trip, we can divide 400 km by 4 h. It wouldn't be constant because we could be driving at different speeds during the trip.

OpenStudy (anonymous):

yea i think your right Auctoratrox, sorry my mistake

OpenStudy (anonymous):

average

OpenStudy (anonymous):

speed ave=d/t=400/4=100 km/h